I need to create a linear scenario and problem. It must have at least two parts to my problem for my classmates to solve. And I have to know the answer to my own question! I'm totally lost! Please help!

Here’s an example below.

Newlyweds just bought a new house to start a family. The original cost of the house is $450,00. The house is depreciated linearly over 10 years, and then have a scrap value of $325,000. Explain the process to write a formula that expresses the house value (v), with reference to age (t) in years. What will be the house value going to be in 5 years? Explain, at what point is the formula useless?
Answer #1
Cost of house $         450,000
Less: Salvage value $         325,000
Depreciable cost $         125,000
Divided by: Useful life $                   10
Annual Depreciation $           12,500
Explain the process to write a formula that expresses the house value (v), with reference to age (t) in years.
House value = 450000 - 12500*Years
v = 450000 - 12500*t
What will be the house value going to be in 5 years?
v = 450000 - 12500*t
v = 450000 - 12500*5
v = 450000 - 62500
v = 387500
at what point is the formula useless?
The house is depreciated linearly over 10 years, Thus, After 10 years is the formula useless.